OpenStudy (help108):

What is the solution to the equation fraction 3 over 5 n minus fraction 4 over 5 equals fraction 1 over 5 n? n = −1 n = 2 n = fraction 5 over 3 n = fraction 5 over 2

alones (alones):

i GAVE YOU the answer in chat.. fine i'll show you the steps >.< 2+y=−3 Simplify both sides of the equation. y+2=−3 then Subtract 2 from both sides. y+2−2=−3−2 y=−5

OpenStudy (help108):

i accidentlly wrote the wrong question do u know the awnser for this one?

alones (alones):

what's the question then -.-

OpenStudy (help108):

i gave u a medal but do u know this anwser?

OpenStudy (help108):

i edited it

OpenStudy (help108):

What is the solution to the equation fraction 3 over 5 n minus fraction 4 over 5 equals fraction 1 over 5 n? n = −1 n = 2 n = fraction 5 over 3 n = fraction 5 over 2

OpenStudy (help108):

What is the solution to the equation fraction 3 over 5 n minus fraction 4 over 5 equals fraction 1 over 5 n? n = −1 n = 2 n = fraction 5 over 3 n = fraction 5 over 2

OpenStudy (help108):

u know the awnser? plezzzz help

alones (alones):

You mean this? 3/5n -4/5 =1/5n 35n−45=15n simplify both sides of the equation. 3/5n+−4/5=15n subtract 1/5n from both sides. 3/5n+−4/5−1/5n=1/5n−1/5n 2/5n+−4/5=0 add 4/5 to both sides. 2/5n+−4/5+4/5=0+4/5 2/5n=4/5 multiply both sides by 5/2. (5/2)*(2/5n)=(5/2)*(4/5) n=2

OpenStudy (help108):

B